Two of the angles of a heptagon are 150 degrees. The other 5 angles are all y. What is the value of y?
Answer by Alan3354(69443)

Two of the angles of a heptagon are 150 degrees. The other 5 angles are all y. What is the value of y?
-------------------
A heptagon has 7 sides. The total of the angles is 180*(n-2) = 900 degrees.
900 - 2*150 = 600
600/5 = 120 degrees = y